.Find the range of the function y = log3 (5 + 4x - x2 )

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
log3 -( x^2-4x+4-4-5 )=log3(9-(x-2)^2)
domain((3-x+2)(3+x-2)>0)
-x+5>0 & x+1>0 -1<x>5
-x+5<0 & x+1<0 no common region
at x=-1 -inf
at x=2 , log3(5+8-4)=log3(9)=2 ,
So the range of function is (-inf,2]

Sher Mohammad
B.Tech, IIT Delhi